# GeoZero: Incentivizing Reasoning from Scratch on Geospatial Scenes

### v1

@@ 6 Overview
  • Section 7: Evaluating whether general MLLMs can perform geospatial reasoning.
  • Section 8: Implementation details of instruction formatting for constructing GeoZero-Raw.
  • Section 9: Detailed data filtering and selection process for producing GeoZero-Hard.
  • Section 10: System prompt design for GeoZero-Hard.
  • Section 11: Calculation process of the thinking-quality score $s_{t}$.
  • Section 12: Calculation of the task-specific answer reward $r_{a}$ for different tasks.
  • Section 13: Specific reinforcement fine-tuning configurations across multiple datasets for GeoZero.
  • Section 14: Further analysis of reasoning behavior.
  • Section 15: Additional qualitative prediction examples of GeoZero.
  • Section 16: Datasheets for GeoZero-Raw.
7 Can General MLLMs Perform Geospatial Reasoning?
@@
To this end, we conduct a diagnostic experiment using the base model of GeoZero, Qwen3-VL-8B-Instruct. To ensure fairness, we align the evaluation configuration with that used for GeoZero in the main paper, where the model is guided by the same system prompt employed in the reinforcement learning stage, specifically the Type1 version without illustrative examples (see Figure 15). We evaluate the model on two remote sensing vision-language datasets, UCM [ucm] and RSVG [rsvg], where the model is instructed to produce an explicit reasoning process before outputting the final answer.
@@
Table 10: Accuracy and thinking activation rate (%) of different models on the UCM and RSVG datasets. w/SP: with the system prompt (Type1) shown in Figure 15. Acc.: accuracy. TAR: thinking activation rate.
Model UCM [ucm] RSVG [rsvg]
@@
GeoZero
(w/ RFT, w/ SP)
95.48
99.05
(96.39% corrected)
(416/420)
50.04
99.43
(1220/1227)
The results indicate that applying system prompts improves the accuracy on both datasets (75.71% $\rightarrow$ 77.14%, 5.95% $\rightarrow$ 6.03%). On UCM, the model performs explicit reasoning on 53.10% of the samples, achieving an accuracy of 80.72% on these samples, which is higher than the overall accuracy of 77.14%. This suggests that general-purpose MLLMs are capable of performing reasoning in certain geospatial contexts, and such reasoning indeed enhances performance.
@@
For the model after supervised fine-tuning (SFT) on the GeoZero-Instruct dataset (notably, without any CoT data), we observe that while the model becomes more familiar with geospatial scenes, leading to improved performance, its reasoning ability is entirely lost. During SFT, the model tends to memorize simple question-answering patterns, which limits its capacity for thinking. This underscores the importance of using reinforcement learning (RL) to activate the geospatial reasoning capabilities of MLLMs.
In contrast, our GeoZero framework, even without any reinforcement fine-tuning (RFT) on individual datasets, consistently produces explicit reasoning traces while improving accuracy. On the UCM dataset, the reasoning activation rate is slightly below 100%, likely because some samples are too simple to require reasoning. On the more complex RSVG dataset, GeoZero demonstrates full activation of geospatial reasoning, reaching a 100% reasoning rate. These findings further validate the rationale behind our research motivation. When applying RFT, both overall performance and the accuracy of reasoned samples improve, despite a slight decrease in the thinking activation rate, possibly due to the model becoming more familiar with the dataset.
@@
8 Instruction Formatting of GeoZero-Raw
We apply instruction formatting to the GeoZero-Raw dataset. To distinguish different tasks, following the conventions of existing remote sensing MLLMs [kuckreja2024geochat, zhan2025skyeyegpt], we prepend task descriptors to user instructions, where [cls], [grounding], [vqa], and [caption] represent scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC) tasks, respectively. Furthermore, to help general-purpose MLLMs better understand the expected output format for different remote sensing vision–language tasks, we introduce task-specific textual hints in addition to the original user instructions. For example, for VG tasks, the instruction is reformulated as: “Please provide the location of the object in the image: {original grounding text}.” To prevent the model from over-relying on these textual hints, we randomly decide whether to include them with a probability of 50%. In addition, to further enhance model generalization, we design more than 20 variations of textual hints for each task, from which one is randomly selected for each sample. The textual hints for different tasks are shown in Figure 10-13, while the converted samples are presented in Figure 14.
9 More Details in Producing GeoZero-Hard
Data Filtering Model. Data filtering model (DFM) is trained via supervised fine-tuning on the GeoZero-Raw dataset, using Qwen2.5-VL-7B-Instruct [qwen25-vl] as the base model. Training is conducted for 3 epochs with a global batch size of 32 across 8 NVIDIA A100 GPUs. We use a learning rate of 1e-4 with a weight decay of 0.1. The vision encoder is kept frozen, while the remaining modules are fine-tuned with LoRA [lora] (rank 8, $\alpha=32$). We adopt bfloat16 precision, FlashAttention [dao2022flashattention], and DeepSpeed [deepspeed] to reduce memory usage and improve training efficiency.
First-Stage Filtering. In the first round of filtering, we run a single forward prediction using DFM on all samples in GeoZero-Raw that belong to the SC, VG, VQA, and IC tasks. Based on the correctness of the predicted output, we determine whether a sample should be retained as a potential hard-sample candidate. The task-specific criteria for judging prediction correctness are defined as follows:
@@ 9 More Details in Producing GeoZero-Hard
10 Prompt Template for GeoZero-Hard
During the reinforcement learning stage, we introduce system prompts to the GeoZero-Hard dataset. To help the model understand the expected answer style for each task, we incorporate a single hypothetical answer example into the system prompt. Similar to the textual hints used in GeoZero-Raw, we randomly decide whether to include these illustrative examples with a probability of 50% to enhance model generalization. We argue that such examples do not introduce human bias, as they are entirely hypothetical and no corresponding images are provided, they merely serve to clarify the expected output format for the model. The specific examples of these system prompts are presented in Figure 15.
11 Thinking-Quality Score $s_{t}$
@@
Equations (6)–(7) jointly ensure that only reasoning processes that are structurally sound, sufficiently informative, semantically diverse, and clearly distinct from the final answer can achieve high thinking-quality scores. The specific computation of $l_{s}$, $p_{r}$, $p_{a}$ and $b_{d}$ are provided as follows:
@@
Semantic Diversity Bonus. For the semantic diversity bonus $b_{d}$, we first split $T$ into $n$ sentences $\{t_{i}\}_{i=1}^{n}$ and obtain their normalized embeddings $\{\mathbf{e}_{i}\}_{i=1}^{n}$ using the bge-small-en-v1.5 model [bge_embedding]. We then quantify sentence-level semantic diversity as
@@
Interestingly, this may also partly explain why our method achieves especially strong improvements over comparison methods on grounding tasks. While gains are observed across all tasks, grounding answers consist of numeric coordinates with essentially no lexical overlap with the reasoning text, and are thus unaffected by this near-uniform penalization.
12 Task-Specific Answer Reward $r_{a}$
@@
where $\cos(f(x),f(y))=\frac{f(x)\cdot f(y)}{\lVert f(x)\rVert\,\lVert f(y)\rVert}.$ In our implementation, $f$ is instantiated by the bge-small-en-v1.5 model [bge_embedding] with normalized embeddings.
@@
Scene Classification
UCM [ucm] 1 5e-5 144
AID [aid] 1 5e-5 144
RESISC-45 [nwpu] 1 1e-4 144
EuroSAT [eurosat] 1 5e-5 144
SIRI-WHU [siri-whu] 1 5e-5 96
WHU-RS19 [whurs19] 1 5e-6 96
Visual Grounding
RSVG [rsvg] 1 5e-5 48
DIOR-RSVG [dior-rsvg] 1 5e-6 96
VRS-VG [li2024vrsbench] 1 5e-6 144
Visual Question Answering
RSVQA-HR-Presence [rsvqa] 1 5e-6 144
RSVQA-HR-Compare [rsvqa] 1 5e-6 144
Image Captioning
UCM-Captions [ucm-captions] 1 5e-6 144
Sydney-Captions [ucm-captions] 3 5e-6 48
NWPU-Captions [nwpu-captions] 1 5e-5 96
RSICD [rsicd] 1 5e-5 144
Challenging Benchmarks
XLRSBench[Wang_2025_CVPR_XLRSBench] 1 5e-6 48
CHOICE[an2025choice] 1 5e-6 48
13 Reinforcement Fine-tuning Settings
@@
14 Further Investigation of Reasoning Behavior
14.1 Overall Characteristics of Model Reasoning.
@@ 14 Further Investigation of Reasoning Behavior
    Figure 6: The distribution of average accuracy and sample frequency on thinking-related metrics, where the statistics are computed from GeoZero’s predictions on the 7500 samples from the DIOR-RSVG [dior-rsvg] test set. The accuracy of samples are represented by IoU score.
    Figure 7: The distribution of thinking-related variables with respect to accuracy, where the statistics are computed from GeoZero’s predictions on 7,500 samples from the DIOR-RSVG [dior-rsvg] test set. The sample accuracy is represented by the IoU score. The intervals are determined based on the quantiles of the IoU distribution, resulting in non-uniform spacing between the bars.
@@
14.2 Reasoning Behavior across Samples.
@@
15 More Examples of Model Predictions
@@ 2. How many instances are there in total (of each type, if appropriate)?
A3: GeoZero-Raw is a curated subset of existing remote sensing vision and vision-language datasets (e.g., VHM-Instruct [pang2025vhm], fMoW [fmow]). It samples diverse ground resolutions, sensor modalities, and scene categories to ensure representative geographic and semantic coverage, while the textual annotations are independently processed and refined by the authors.
@@ 10. Is the dataset self-contained, or does it link to or otherwise rely on external resources?
A10: The dataset is self-contained in terms of structure and usability. All necessary textual annotations required for training are included. The image data originates from publicly available remote sensing datasets (e.g., VHM-Instruct [pang2025vhm], fMoW [fmow]). Users can access the images via the provided links.
@@ 3. If the dataset is a sample from a larger set, what was the sampling strategy?
A3: The dataset is initially built upon the VHM-Instruct dataset [pang2025vhm]. Since we mainly focus on scene classification, visual grounding, visual question answering, and image captioning tasks, we further supplement additional samples from the training sets of existing datasets for each task. This process alleviates the task imbalance issue in VHM-Instruct and ensures comprehensive task coverage across different modalities.
@@ 5. Over what timeframe was the data collected?
16.4 Preprocessing/Cleaning/Labeling

### v2

@@ Appendix A Overview
  • Section B: Evaluating whether general MLLMs can perform geospatial reasoning.
  • Section C: Implementation details of instruction formatting for constructing GeoZero-Raw.
  • Section D: Detailed data filtering and selection process for producing GeoZero-Hard.
  • Section E: System prompt design for GeoZero-Hard.
  • Section F: Calculation process of the thinking-quality score $s_{t}$.
  • Section G: Calculation of the task-specific answer reward $r_{a}$ for different tasks.
  • Section H: Specific reinforcement fine-tuning configurations across multiple datasets for GeoZero.
  • Section I: Further analysis of reasoning behavior.
  • Section J: Additional qualitative prediction examples of GeoZero.
  • Section K: Datasheets for GeoZero-Raw.
Appendix B Can General MLLMs Perform Geospatial Reasoning?
@@
To this end, we conduct a diagnostic experiment using the base model of GeoZero, Qwen3-VL-8B-Instruct. To ensure fairness, we align the evaluation configuration with that used for GeoZero in the main paper, where the model is guided by the same system prompt employed in the reinforcement learning stage, specifically the Type1 version without illustrative examples (see Figure 8). We evaluate the model on two remote sensing vision-language datasets, UCM [65] and RSVG [55], where the model is instructed to produce an explicit reasoning process before outputting the final answer.
@@
Table 9: Accuracy and thinking activation rate (%) of different models on the UCM and RSVG datasets. GeoZero uses the system prompt (SP) by default. w/SP: with the system prompt (Type1) shown in Figure 8. Acc.: accuracy. TAR: thinking activation rate.
Model UCM [65] RSVG [55]
@@
The results indicate that directly applying system prompts cannot improve the accuracy of general MLLM on both datasets (75.71% $\rightarrow$ 77.14%, 26.41% $\rightarrow$ 24.04%). On UCM, the model performs explicit reasoning on 53.10% of the samples, achieving an accuracy of 80.72% on these samples, which is higher than the overall accuracy of 77.14%. This suggests that general-purpose MLLMs are capable of performing reasoning in certain geospatial contexts, and such reasoning indeed enhances performance.
@@
In contrast, our GeoZero framework, consistently produces explicit reasoning traces while improving accuracy. On the UCM dataset, the reasoning activation rate is slightly below 100%, likely because some samples are too simple to require reasoning. On the more complex RSVG dataset, GeoZero demonstrates full activation of geospatial reasoning, reaching a 100% reasoning rate. These findings further validate the rationale behind our research motivation.
@@
Appendix C Instruction Formatting of GeoZero-Raw
We apply instruction formatting to the GeoZero-Raw dataset. To distinguish different tasks, following the conventions of existing remote sensing MLLMs [28, 68], we prepend task descriptors to user instructions, where [cls], [grounding], [vqa], and [caption] represent scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC) tasks, respectively. For samples that do not belong to these four vision–language tasks, no task descriptors are added. Furthermore, to help general-purpose MLLMs better understand the expected output format for different remote sensing vision–language tasks, we introduce task-specific textual hints in addition to the original user instructions.. For example, for VG tasks, the instruction is reformulated as: “Please provide the location of the object in the image: {original grounding text}.” To prevent the model from over-relying on these textual hints, we randomly decide whether to include them with a probability of 50%. In addition, to further enhance model generalization, we design more than 20 variations of textual hints for each task, from which one is randomly selected for each sample. The textual hints for different tasks are shown in Figure 11-14, while the converted samples are presented in Figure 15.
Appendix D More Details in Producing GeoZero-Hard
Data Splitting. To reduce potential bias and mitigate trivial memorization effects, we first randomly sample approximately 300K instances from GeoZero-Raw to form a validation subset, denoted as GeoZero-Raw-val, which is reserved exclusively for hard-sample selection. The remaining samples constitute GeoZero-Raw-train and are used solely for training.
This split ensures that the difficulty assessment is performed on samples unseen during subsequent training.
Data Filtering Model. The data filtering model (DFM) is trained via supervised fine-tuning on the GeoZero-Raw-train dataset, using Qwen2.5-VL-7B-Instruct [5] as the base model. Training is conducted for 3 epochs with a global batch size of 32 across 8 NVIDIA A100 GPUs. We use a learning rate of 1e-4 with a weight decay of 0.1. The vision encoder is kept frozen, while the remaining modules are fine-tuned with LoRA [20] (rank 8, $\alpha=32$). We adopt bfloat16 precision, FlashAttention [12], and DeepSpeed [49] to reduce memory usage and improve training efficiency.
First-Stage Filtering. In the first stage, we apply the trained DFM to all samples in GeoZero-Raw-val that belong to the SC, VG, VQA, and IC tasks, and perform a single forward inference pass.
Samples that are predicted incorrectly are retained for further consideration. Here, we emphasize that misprediction is not assumed to be a definitive indicator of reasoning complexity; rather, it serves as a practical proxy for identifying samples that are challenging for a strong instruction-tuned MLLM.
The task-specific criteria for judging prediction correctness are defined as follows:
@@ Appendix D More Details in Producing GeoZero-Hard
Appendix E Prompt Template for GeoZero-Hard
During the RL stage, we introduce system prompts to the GeoZero-Hard dataset. To help the model understand the expected answer style for each task, we incorporate a single hypothetical answer example into the system prompt. Similar to the textual hints used in GeoZero-Raw, we randomly decide whether to include these illustrative examples with a probability of 50% to enhance model generalization. We argue that such examples do not introduce human bias, as they are entirely hypothetical and no corresponding images are provided, they merely serve to clarify the expected output format for the model. The specific examples of these system prompts are presented in Figure 8.
Appendix F Thinking-Quality Score $s_{t}$
@@
Equations (6)-(7) jointly ensure that only reasoning processes that are structurally sound, sufficiently informative, semantically diverse, and clearly distinct from the final answer can achieve high thinking-quality scores. The specific computation of $l_{s}$, $p_{r}$, $p_{a}$ and $b_{d}$ are provided as follows:
@@
Semantic Diversity Bonus. For the semantic diversity bonus $b_{d}$, we first split $T$ into $n$ sentences $\{t_{i}\}_{i=1}^{n}$ and obtain their normalized embeddings $\{\mathbf{e}_{i}\}_{i=1}^{n}$ using the bge-small-en-v1.5 model [63]. We then quantify sentence-level semantic diversity as
@@
Interestingly, this may also partly explain why our method achieves especially strong improvements over comparison methods on grounding tasks. While gains are observed across all tasks, grounding answers consist of numeric coordinates with essentially no lexical overlap with the reasoning text. As a result, this near-uniform penalization has limited influence on the thinking trajectories in VG tasks.
Appendix G Task-Specific Answer Reward $r_{a}$
@@
where $\cos(f(x),f(y))=\frac{f(x)\cdot f(y)}{\lVert f(x)\rVert\,\lVert f(y)\rVert}.$ In our implementation, $f$ is instantiated by the bge-small-en-v1.5 model [63] with normalized embeddings.
@@
UCM [65] 1 5e-5 144
RSVG [55] 1 5e-5 48
Appendix H Reinforcement Fine-tuning Settings
@@
Appendix I Further Investigation of Reasoning Behavior
I.1 Overall Characteristics of Model Reasoning.
@@ Appendix I Further Investigation of Reasoning Behavior
    Figure 9: The distribution of average accuracy and sample frequency on thinking-related metrics, where the statistics are computed from GeoZero’s predictions on the 7500 samples from the DIOR-RSVG [67] test set. The accuracy of samples are represented by IoU score.
    Figure 10: The distribution of thinking-related variables with respect to accuracy, where the statistics are computed from GeoZero’s predictions on 7,500 samples from the DIOR-RSVG [67] test set. The sample accuracy is represented by the IoU score. The intervals are determined based on the quantiles of the IoU distribution, resulting in non-uniform spacing between the bars.
@@
I.2 Reasoning Behavior across Samples.
@@
Appendix J More Examples of Model Predictions
@@ 2. How many instances are there in total (of each type, if appropriate)?
A3: GeoZero-Raw is a curated subset of existing remote sensing vision and vision-language datasets (e.g., VHM-Instruct [47], fMoW [9]). It samples diverse ground resolutions, sensor modalities, and scene categories to ensure representative geographic and semantic coverage, while the textual annotations are independently processed and refined by the authors.
@@ 10. Is the dataset self-contained, or does it link to or otherwise rely on external resources?
A10: The dataset is self-contained in terms of structure and usability. All necessary textual annotations required for training are included. The image data originates from publicly available remote sensing datasets (e.g., VHM-Instruct [47], fMoW [9]). Users can access the images via the provided links.
@@ 3. If the dataset is a sample from a larger set, what was the sampling strategy?
A3: The dataset is initially built upon the VHM-Instruct dataset [47]. Since we mainly focus on scene classification, visual grounding, visual question answering, and image captioning tasks, we further supplement additional samples from the training sets of existing datasets for each task. This process alleviates the task imbalance issue in VHM-Instruct and ensures comprehensive task coverage across different modalities.
@@ 5. Over what timeframe was the data collected?
K.4 Preprocessing/Cleaning/Labeling